# AI-Slop to AI-Polish? Aligning Language Models through Edit-Based Writing Rewards and Test-time Computation

### v2

@@ 1 Introduction
Despite LLMs showing impressive performance in math and coding, their ability to write high-quality text has been rather pedestrian. Recent work from Chakrabarty et al. (2024b) shows how text generated from widely used LLMs are often rife with clichés, purple prose, poor sentence structure, and unnecessary exposition. This stems from several challenges. Unlike math or coding, writing lacks verifiable rewards. While it would be possible to train a model to write better text by having humans label examples of “good” and “bad” writing, it is challenging due to the required expertise. Self-evaluation using LLMs has proven useful in reward modeling and constitutional AI (Bai et al., 2022), but relying on uncalibrated humans or LLMs for feedback (Lee et al., 2023; Gao et al., 2024) on subjective tasks like writing can lead to reward hacking (Pan et al., 2024) and alignment issues. Recent work from Panickssery et al. (2024) shows the self-aggrandizing nature of LLMs, as evidenced in Table 1 where they prefer their own writing over Nobel Prize winners’ work.
@@
The surge in AI writing assistance demands urgent alignment of AI-generated text with human preferences. Recent work from Gooding et al. (2025) show how LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. They highlight the need for models to better assess the quality and impact of suggested actions, both during generation and across multi-step refinement. Binary preference feedback between paired examples is the most common alignment method for LLMs (Christiano et al., 2017), but it has a significant drawback. The paired outputs may differ in several ways and could be equally worse in terms of quality (Casper et al., 2023; Lambert & Calandra, 2023). [^1] Forcing annotators to choose between two undesirable outputs doesn’t improve alignment. In the current design of RLHF, annotators are not allowed to pick neither Recent work from Chakrabarty et al. (2024b) shows how identifying and editing problematic response segments effectively improves AI alignment. This also reflects the Reviewing phase in the cognitive process model of writing (Hayes et al., 1987), where humans evaluate and revise text. They release LAMP (Language model Authored, Manually Polished), a corpus of 1282 $<AI-generated,Expert-Edited>$ pairs with implicit preference ($edited>original\_draft$) to improve AI writing (see Table 4 in Appendix B). Additionally, each paragraph pair includes normalized scores (1-10) reflecting writing quality before and after editing.
Which is better quality writing?
@@ 3 Writing Quality Reward Models
We rely on the LAMP (Language model Authored, Manually Polished) corpus from Chakrabarty et al. (2024b) to train reward models. As illustrated in Figure 2, each sample in LAMP consists of a writing instruction and two paragraphs that match this instruction. The paragraphs in LAMP range from 150 to 400 words, and span across fiction and non-fiction. Table 4 in Appendix B shows a sample from LAMP, highlighting the edits implemented by an expert to improve writing quality. We use three methods to transform LAMP samples into training and validation data points for our models: pairwise (P), scalar (R), and combined (PR). With the P method, each data point presents two paragraphs as input (1 and 2) and requires a binary classification output indicating which paragraph has higher writing quality (i.e., the output is 1 or 2). Each LAMP sample is duplicated into two P data points by considering both paragraph orders (AI-generated, Expert-Edited $\to$ 2) and (Expert-Edited, AI-generated $\to$ 1). With the R method, each data point takes a single paragraph as input and outputs a regression value predicting the quality score of that paragraph. Since each LAMP sample contains two paragraphs (before and after edit), it generates two R data points. The PR method combines both approaches, yielding four data points per LAMP sample (two from P and two from R). There are a total of 1,282 samples in LAMP, and we follow the author’s split divisions of 1,000 training, 67 validation, and 215 test samples. Applying the data transformation described above, the P, R, and PR variants of the training data we obtain consist of 2,000, 2,000, and 4,000 training data points, respectively. For our experiments, we trained both generative LLMs (Llama3.1 (Dubey et al., 2024)) and encoder-only models (ModernBert (Warner et al., 2024)).
@@
We follow the standard approach introduced in the original BERT paper (Devlin et al., 2019) to add and finetune two task-specific heads to a ModernBERT-Large model (Warner et al., 2024). The input data points contain either one paragraph (for R data points) or two paragraphs (for P data points), which are encoded jointly with a pre-defined separator token when needed. For each paragraph, we compute a “paragraph vector” by pooling the last layer’s activations across all tokens in that paragraph. These paragraph vectors serve as input to either a regression (R) or classification (P) head. The regression head transforms the vector through a learned linear projection from the model’s inner dimension to a scalar, followed by a scaled sigmoid to align with the 1-10 score range. The classification head is aparametric, using a cosine similarity operation between the two paragraph vectors. We use mean-squared error loss for R data points and cross entropy for P data points. Following convention for encoder-only models, we finetune the entire model’s weights (Devlin et al., 2019). We selected ModernBERT-Large, the largest available model, for our experiments. We fine-tuned three variants: MBERT-WQRM-P, MBERT-WQRM-R, and MBERT-WQRM-PR, each on their corresponding data variants. Hyperparameters, including learning rate and number of epochs, were optimized by minimizing validation loss.
@@
We use a standard approach to finetune generative transformer architectures by transforming the classification and regression tasks into sequence-to-sequence problems. Specifically, we adopt a JSON format on the output side, as shown in the Appendix’s Table 5. We follow convention and use the QLora (Dettmers et al., 2023) parameter-efficient tuning method with FSDP (Zhao et al., 2023), and use a cross-entropy loss.
A key advantage of generative methods is their potential to produce natural-language rationales alongside their predictions, which can be leveraged for the interpretability of the model responses. Prior work from Wiegreffe et al. (2020) has demonstrated label-rationale association as an important property for faithfulness of model response. Additionally researchers (Ludan et al., 2023; Hase & Bansal, 2021) have argued for incorporating explanations in both model input or output as way to make it robust against spurious cues. Since the original LAMP corpus did not collect rationales from experts during annotation, we introduce a procedure to augment LAMP with LLM-generated silver rationales. Specifically, we first collected five examples that either contrast one paragraph’s strength over the other (P-style) or holistically critique or praise the paragraph (R-style) from two professional writers. Writers were specifically instructed to cite excerpts from the paragraphs to justify their rationales. We then use these expert-written rationales as demonstration examples in an in-context instruction and generate rationales using the Claude3.5 Sonnet model  [^3] Considered a top-performing model for writing tasks at the time of experiments.. The Appendix’s Table 6 in Appendix D shows examples of rationale generated by Claude3.5 Sonnet.
@@ 4 The Writing Quality Benchmark
To assess writing quality prior work has argued for evaluation by professionals (ones with writing experience). Nevertheless, some writing quality preference datasets are based on crowd-sourced judgments. We include four datasets based on expert judgments and one dataset based on crowd-sourced annotation (LM Arena) to represent both perspectives in the benchmark. Finally, we selected two datasets with long responses (Art or Artifice, LM Arena) and three with shorter responses ranging from 200-400 words, ensuring that models that perform well on the benchmark are capable of judging writing quality irrespective of length. Appendix D.1 details the procedure we followed to extract and standardize each dataset.
@@ 4.1 Experimental Results on WQ
Our experiments on the WQ benchmark include four classes of models. First, Zero-Shot (ZS) and Few-Shot (FS) methods with top-performing instruction-tuned LLMs. We included both non-reasoning (GPT-4o) and reasoning models (Deepseek-R1, O1). Second, a top-performing generic reward model – SkyWork-8b-v0.2 – based on results on the RewardBench leaderboard (Lambert et al., 2024). Third, we include the Pangram AI-detector  [^4] https://www.pangram.com/dashboard?type=text, accessed through API. Finally, the trained WQRM models in generative and encoder-only settings as described in Section 3. Models that can produce pairwise judgments (such as SkyWork or WQRM-P models) were used as is, but for models that produce scalar rewards (WQRM-R, Pangram), a scalar reward was computed for each response, and inequality was applied to emit a pairwise preference (resolving ties randomly).
@@ 5.1 Generating edits via Supervised Finetuning
Prior work from Chakrabarty et al. (2024b) shows experimentally that LLMs’ text idiosyncrasies (cliches, redundancy, lack of subtext, etc.) can be mitigated through self-editing in an in-context setup. Borrowing motivation from them we teach LLMs how to improve their response via edits. Figure 3 illustrates the three components of the editing pipeline. Given a first draft response to an instruction from any given LLM, the first step consists of identifying and listing idiosyncrasies: spans in the first draft that can be rephrased to improve overall writing quality. For each identified idiosyncrasy, a second stage consists in rewriting the idiosyncrasy. This is framed as an executable edit (Laban et al., 2023), where each edit consists of replacing an original string in a draft with an improved version. The third step simply executes all edits (by applying a series of string replace operations) to obtain the final edited draft. While Chakrabarty et al. (2024b) implemented this through prompt-chaining (Wu et al., 2022) with few-shot examples, we improved efficiency by supervised fine-tuning of GPT-4o and Llama3.1 70B based on the entire LAMP training set. The training input consists of the first draft alongside the entire edit interaction trace (detect, rewrite, execute) in a step-by-step chain of thought prompt, and the output is the edited paragraph. See Appendix F for an example COT prompt.
@@ 5.2 Selecting edited response by leveraging Test-Time Compute
Recent work from Snell et al. (2024) shows that test-time compute can be scaled optimally by using a reward model to search over the space of solutions. This approach typically involves generating multiple candidate responses and using a verifier to select an optimal response (Cobbe et al., 2021). The most popular technique to increase test-time compute is Best-of-N sampling also known as Rejection Sampling, in which N candidates are generated independently. The reward model is then used to score each candidate, and the top-scoring candidate is selected. While test-time scaling is effective for reasoning tasks, our work aims to measure whether it is a practical strategy to improve human-AI alignment in subjective tasks such as writing.
Next we describe the validation study with experts to measure how well calibrated our WQRMs are to human judgment and whether additional test-time computation leads to meaningful improvements in AI writing quality.
@@ 6.1 Study Findings
Figure 4: Results and analysis of WQRM based: (4(a)) distribution of preference based on 300 expert triplet rankings, (4(b)) calibration between gap in WQRM scores and matching expert preference, and (4(c)) applying experts edits gradually to a draft leads to gradual reward gains.
Figure 4 summarizes findings from the expert annotation. In Figure 4(a), we plot the distribution of rankings across all triplets. Best Edit candidates were most preferred overall with an average rank of 1.58, followed by random edit (2.09) and first draft (2.26). The breakdown of rankings across domains (fiction, non-fiction, marketing) or LLM (GPT-4o vs. Llama 3.1) is presented in Appendix G. In short, Best Edit achieves the top rank in all conditions, confirming the generalization of WQRM scores across conditions.
If the reward model is well-calibrated, the gap in WQRM score between two responses should be indicative of the qualitative gap between the responses. For example, if responses A and B receive a score of 4 and 6, the difference in quality between the pair should be larger than if the scores were 4 and 4.5. To inspect the calibration of WQRM, we compute the gap in WRQM between all pairs of annotated responses, and plot the relationship between the gap and the level of agreement with the expert annotations. As shown in Figure 4(b), WQRM gap correlates positively in terms of agreement with experts: when pair of responses differs in WQRM score by 0.5 points or less, an individual expert is only 55% likely to prefer the higher-scoring response. When the gap is above 3.0, the proportion goes up to 80%, indicating much stronger agreement. Besides agreement with individual experts, we plot agreement levels with the majority rank based on three expert annotations (green line), finding stronger positive agreement than with individual experts. In short, we find evidence that WQRM is well-calibrated: a wider gap in scores between two responses is evidence that an expert (or group of experts) would be more likely to prefer the higher-scoring response over the lower-scoring response.
@@ 7 How does content affect writing quality?
Figure 5: Writing quality analysis of human-written and LLM-generated texts according to the WQRM model on (5(a)) less and (5(b)) more content detail in the writing prompt. Prompts with less content detail average 30 words, whereas prompts with more content detail average 180.
@@
Effectively judging writing quality has a broader impact on both understanding and improving LLM writing. Writing quality, however, is often closely tied to content. It is widely known that LLMs face challenges in coming up with novel ideas (content planning), which can make their writing appear trite or hackneyed. Conversely, even when provided with detailed original content, they might struggle to stitch everything together while ensuring that it conforms to good writing standards (i.e., fewer clichés, avoiding revealed subtext, and minimal use of ornamental/purple prose).To better understand how content effects writing quality, we conducted an analysis involving several LLMs on how they write with or without detailed content in the writing prompt. We leverage the 50 writing instructions in Style Mimic data, and convert them into two variants: a writing prompt with less content detail that spans 30 words (e.g., “A family Christmas unfolds through emotional reflections on a father’s new family, a daughter’s excuse to stay behind, and the complex dynamics of grief and blended identities.”) and a writing prompt with more content detail that spans 150-200 words long (see Table 10 in Appendix). For each writing prompt, Style Mimic provides an original excerpt from an award-winning author and a second excerpt written by an MFA student attempting to mimic the original author’s style. Each sample in the Style Mimic data comes with the detailed content used for 5(b).
Since WQRM was only trained on samples from LAMP, which consists of AI-generated paragraphs edited by MFA students, it doesn’t fully know how to reward higher-quality human writing. For this purpose, we added 100 paragraphs written by 5 award-winning authors (20 each) to our training data. We chose 5 authors who were part of the Style Mimic data. Each paragraph written by an award-winning author was assigned a score of 10.0. Even within writing from trained professionals, there is significant variability. To address this we source an additional 80 independent paragraphs written by MFA students published in prestigious literary magazines such as Electric Lit, Joyland, Paris Review and add to our training data. Each paragraph written by an MFA student was assigned a score of 7.5  [^7] This was a design decision where 5 is average and 10 is the best, so we wanted to choose the mean of both ends. Publication at a venue already means these paragraphs have undergone scrutiny and are of decent quality. After adding these 180 samples to LAMP-PR training set, we retrained WQRM.
@@
Figure 5(a) shows the writing quality of several LLMs judged by the new WQRM model when detailed content is not provided in the writing prompt. Award-winning authors score higher, with a median score of 8.9. Writing quality distributions for LLMs differ broadly from award-winning authors. All models have much higher variance than the award winning authors and achieve a median score of 4.8 to 6.6. Since WQRM was only trained on AI-generated paragraphs edited by MFA students and data from only 5 award-winning authors, the ability of the model to score 50 author-written texts higher than all LLMs provides evidence of generalization from the model. It should also be noted that our WQRMs don’t find GPT4.5, otherwise touted as the best LLM for writing, to be of any better quality. The significant gap between award-winning authors and LLMs shows that in the absence of original good-quality content, all LLMs are poor writers.

### v3

@@ 1 Introduction
Despite LLMs showing impressive performance in math and coding, their ability to write high-quality text has been rather pedestrian. Recent work from Chakrabarty et al. (2024b) shows how text generated from widely used LLMs are often rife with clichés, purple prose, poor sentence structure, and unnecessary exposition. This stems from several challenges. Unlike math or coding, writing lacks verifiable rewards. While it would be possible to train a model to write better text by having humans label examples of “good” and “bad” writing, it is challenging due to the required expertise. Self-evaluation using LLMs has proven useful in reward modeling and constitutional AI (Bai et al., 2022), but relying on uncalibrated humans or LLMs for feedback (Lee et al., 2023; Gao et al., 2024) on subjective tasks like writing can lead to reward hacking (Pan et al., 2024) and alignment issues. Recent work from Panickssery et al. (2024) shows the self-aggrandizing nature of LLMs, as evidenced in Table 3 where they prefer their own writing over Nobel Prize winners’ work. For the purpose of this paper we define good writing quality as writing that doesn’t contain disproportionate amount of peculiar words or phrases, has fewer cliches or hackneyed expressions, is not unnecessarily ornamental as well as doesn’t have a overly saccharine and polished tone or voice.
@@
The surge in AI writing assistance demands urgent alignment of AI-generated text with human preferences. Recent work from Gooding et al. (2025) show how LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. They highlight the need for models to better assess the quality and impact of suggested actions, both during generation and across multi-step refinement. Binary preference feedback between paired examples is the most common alignment method for LLMs (Christiano et al., 2017), but it has a significant drawback. The paired outputs may differ in several ways and could be equally worse in terms of quality (Casper et al., 2023; Lambert & Calandra, 2023). [^1] Forcing annotators to choose between two undesirable outputs doesn’t improve alignment. In the current design of RLHF, annotators are not allowed to pick neither Recent work from Chakrabarty et al. (2024b) shows how identifying and editing problematic response segments effectively improves AI alignment. This also reflects the Reviewing phase in the cognitive process model of writing (Hayes et al., 1987), where humans evaluate and revise text. They release LAMP (Language model Authored, Manually Polished), a corpus of 1282 <AIgenerated,ExpertEdited><AI-generated,Expert-Edited> pairs with implicit preference (edited>original_draftedited>original\_draft) to improve AI writing (see Table 4 in Appendix A.1). Additionally, each paragraph pair includes normalized scores (1-10) reflecting writing quality before and after editing.
@@ 3 Writing Quality Reward Models
We rely on the LAMP (Language model Authored, Manually Polished) corpus from Chakrabarty et al. (2024b) to train reward models. As illustrated in Figure 2, each sample in LAMP consists of a writing instruction and two paragraphs that match this instruction. The paragraphs in LAMP range from 150 to 400 words, and span across fiction and non-fiction. Table 4 in Appendix A.1 shows a sample from LAMP, highlighting the edits implemented by an expert to improve writing quality. We use three methods to transform LAMP samples into training and validation data points for our models: pairwise (P), scalar (R), and combined (PR). With the P method, each data point presents two paragraphs as input (1 and 2) and requires a binary classification output indicating which paragraph has higher writing quality (i.e., the output is 1 or 2). Each LAMP sample is duplicated into two P data points by considering both paragraph orders (AI-generated, Expert-Edited $\to$ 2) and (Expert-Edited, AI-generated $\to$ 1). With the R method, each data point takes a single paragraph as input and outputs a regression value predicting the quality score of that paragraph. Since each LAMP sample contains two paragraphs (before and after edit), it generates two R data points. The PR method combines both approaches, yielding four data points per LAMP sample (two from P and two from R). There are a total of 1,282 samples in LAMP, and we follow the author’s split divisions of 1,000 training, 67 validation, and 215 test samples. Applying the data transformation described above, the P, R, and PR variants of the training data we obtain consist of 2,000, 2,000, and 4,000 training data points, respectively. For our experiments, we trained both generative LLMs (Llama3.1 (Dubey et al., 2024)) and encoder-only models (ModernBert (Warner et al., 2024)).
@@
We follow the standard approach introduced in the original BERT paper (Devlin et al., 2019) to add and finetune two task-specific heads to a ModernBERT-Large model (Warner et al., 2024). The input data points contain either one paragraph (for R data points) or two paragraphs (for P data points), which are encoded jointly with a pre-defined separator token when needed. For each paragraph, we compute a “paragraph vector” by pooling the last layer’s activations across all tokens in that paragraph. These paragraph vectors serve as input to either a regression (R) or classification (P) head. The regression head transforms the vector through a learned linear projection from the model’s inner dimension to a scalar, followed by a scaled sigmoid to align with the 1-10 score range. The classification head is aparametric, using a cosine similarity operation between the two paragraph vectors. We use mean-squared error loss for R data points and cross entropy for P data points. Following convention for encoder-only models, we finetune the entire model’s weights (Devlin et al., 2019). We selected ModernBERT-Large, the largest available model, for our experiments. We fine-tuned three variants: MBERT-WQRM-P, MBERT-WQRM-R, and MBERT-WQRM-PR, each on their corresponding data variants. Hyperparameters, including learning rate and number of epochs, were optimized by minimizing validation loss. PR models can be used in either P- or R-mode at test-time. Initial evaluation indicated that PR models achieve higher performance in R-mode, and as such we used all PR models in R-mode by default during evaluation.
@@
We finetune generative transformer architectures by converting classification and regression tasks to sequence-to-sequence problems using JSON output format (Table 5). We employ QLora (Dettmers et al., 2023) parameter-efficient tuning with FSDP (Zhao et al., 2023) and cross-entropy loss. Generative methods can produce natural-language rationales alongside predictions for interpretability. Wiegreffe et al. (2020) demonstrated label-rationale association as essential for response faithfulness, while (Ludan et al., 2023; Hase & Bansal, 2021) argued for incorporating explanations in model input/output to improve robustness against spurious cues. Since LAMP lacks expert rationales, we augment it with LLM-generated silver rationales. We collected five examples from professional writers showing either paragraph strength contrasts (P-style) or holistic critiques/praise (R-style), instructing them to cite specific excerpts. These expert rationales serve as demonstrations for Claude3.5 Sonnet [^3] Considered a top-performing model for writing tasks at the time of experiments. to generate rationales (examples in Table 6, Appendix A.3).
@@ 4 The Writing Quality Benchmark
To assess writing quality prior work has argued for evaluation by professionals (ones with writing experience). Nevertheless, some writing quality preference datasets are based on crowd-sourced judgments. We include four datasets based on expert judgments and one dataset based on crowd-sourced annotation (LM Arena) to represent both perspectives in the benchmark. Finally, we selected two datasets with long responses (Art or Artifice, LM Arena) and three with shorter responses ranging from 200-400 words, ensuring that models that perform well on the benchmark are capable of judging writing quality irrespective of length. Appendix A.4 details the procedure we followed to extract and standardize each dataset. Appendix A.5 provides an analysis we conducted on the relative difficulty of each dataset in the benchmark, finding that the five selected datasets provide a breadth of coverage in terms of difficulty.
@@ 4.1 Experimental Results on WQ
Our experiments on the WQ benchmark include four classes of models. First, Zero-Shot (ZS) and Few-Shot (FS) methods with top-performing instruction-tuned LLMs. We included both non-reasoning (GPT-4o) and reasoning models (Deepseek-R1, O1). Second, a top-performing generic reward model – SkyWork-8b-v0.2 – based on results on the RewardBench leaderboard (Lambert et al., 2024). Third, we include the Pangram AI-detector  [^4] https://www.pangram.com/dashboard?type=text, accessed through API. Finally, the trained WQRM models in generative and encoder-only settings as described in Section 3. Models that can produce pairwise judgments (such as SkyWork or WQRM-P models) were used as is, but for models that produce scalar rewards (WQRM-R, Pangram), a scalar reward was computed for each response, and inequality was applied to emit a pairwise preference. Scalar rewards can theoretically lead to a tie (a score difference of less than an epsilon like 0.001), but we observe few of these in practice (less than 0.1% of pairs), and resolve those randomly.
@@
We observe that generative WQRM models perform best in P-mode, whereas encoder models perform best in R-mode. We emit a hypothesis for this reversal of relationship, related to the choice of loss. The generative models (Llama) are trained with a sequence-to-sequence loss, whereas the encoder-only models (MBert) are trained with custom losses (pairwise classification for P, mean-squared error for R). In other words, LLama training on the reward-based data is more similar to 10-way classification than actual score regression, whereas the MBert training makes better use of the reward-based data. This leads the MBERT-R models to outperform MBert-P models, whereas the reverse is true for the LLama models, as they are not able to properly take advantage of the R-based data.
@@ 5.1 Generating edits via Supervised Finetuning
Prior work from Chakrabarty et al. (2024b) shows experimentally that LLMs’ text idiosyncrasies (cliches, redundancy, lack of subtext, etc.) can be mitigated through self-editing in an in-context setup. Borrowing motivation from them we teach LLMs how to improve their response via edits. Figure 6 illustrates the three components of the editing pipeline. Given a first draft response to an instruction from any given LLM, the first step consists of identifying and listing idiosyncrasies: spans in the first draft that can be rephrased to improve overall writing quality. For each identified idiosyncrasy, a second stage consists in rewriting the idiosyncrasy. This is framed as an executable edit (Laban et al., 2023), where each edit consists of replacing an original string in a draft with an improved version. The third step simply executes all edits (by applying a series of string replace operations) to obtain the final edited draft. While Chakrabarty et al. (2024b) implemented this through prompt-chaining (Wu et al., 2022) with few-shot examples, we improved efficiency by supervised fine-tuning of GPT-4o and Llama3.1 70B based on the entire LAMP training set. The training input consists of the first draft alongside the entire edit interaction trace (detect, rewrite, execute) in a step-by-step chain of thought prompt, and the output is the edited paragraph. See Appendix A.7 for an example COT prompt.
@@ 5.2 Selecting edited response by leveraging Test-Time Compute
Recent work from Snell et al. (2024) shows that test-time compute can be scaled optimally by using a reward model to search over the space of solutions. This approach typically involves generating multiple candidate responses and using a verifier to select an optimal response (Cobbe et al., 2021). The most popular technique to increase test-time compute is Best-of-N sampling also known as Rejection Sampling, in which N candidates are generated independently. The reward model is then used to score each candidate, and the top-scoring candidate is selected. While test-time scaling is effective for reasoning tasks, our work aims to measure whether it is a practical strategy to improve human-AI alignment in subjective tasks such as writing.Next we describe the validation study with experts to measure how well calibrated our WQRMs are to human judgment and whether additional test-time computation leads to meaningful improvements in AI writing quality.
@@ 6.1 Study Findings
Figure 3: Results and analysis of WQRM based: (3(a)) distribution of preference based on 300 expert triplet rankings, (3(b)) calibration between gap in WQRM scores and matching expert preference, and (3(c)) applying experts edits gradually to a draft leads to gradual reward gains.
Figure 3 summarizes findings from the expert annotation. In Figure 3(a), we plot the distribution of rankings across all triplets. Best Edit candidates were most preferred overall with an average rank of 1.58, followed by random edit (2.09) and first draft (2.26). The breakdown of rankings across domains (fiction, non-fiction, marketing) or LLM (GPT-4o vs. Llama 3.1) is presented in Appendix A.8. In short, Best Edit achieves the top rank in all conditions, confirming the generalization of WQRM scores across conditions.
If the reward model is well-calibrated, the WQRM score gap between responses should indicate their qualitative difference. For example, responses scoring 4 and 6 should have a larger quality gap than those scoring 4 and 4.5. To inspect WQRM calibration, we computed the WQRM gap between all annotated response pairs and plotted it against expert annotation agreement. As shown in Figure  3(b), WQRM gap positively correlates with expert agreement: when responses differ by $\leq 0.5$ points, individual experts prefer the higher-scoring response only 55% of the time. When the gap exceeds 3.0, this increases to 80%. Agreement with majority rank based on three expert annotations (green line) shows even stronger positive correlation. In short, we find evidence that WQRM is well-calibrated: a wider gap in scores between two responses is evidence that an expert (or group of experts) would be more likely to prefer the higher-scoring response over the lower-scoring response.
@@ 7 How does content affect writing quality?
Figure 4: Writing quality analysis of human-written and LLM-generated texts according to WQRM on (4(a)) less and (4(b)) more content detail in the writing prompt. Prompts with less content detail average 30 words, whereas prompts with more content detail average 180.
@@
Effectively judging writing quality impacts both understanding and improving LLM writing. Writing quality is however closely tied to content. Its known that LLMs struggle with novel ideas (content planning), making their writing appear trite. Even with detailed original content, they struggle to maintain good writing standards (avoiding clichés, revealing subtext, and introducing purple prose). To understand how content affects writing quality, we analyzed writing from several LLMs with and without detailed content. We used 50 writing instructions from Style Mimic data, creating two variants: a 30-word prompt with less detail (e.g., ”A family Christmas unfolds through emotional reflections on a father’s new family, a daughter’s excuse to stay behind, and the complex dynamics of grief and blended identities.”) and a 150-200 word detailed prompt (Table 10 in Appendix). Style Mimic provides an original excerpt from an award-winning author and an MFA student’s attempt to mimic that style for each prompt. Each sample includes the detailed content used for 4(b).
@@
Since WQRM was only trained on samples from LAMP, which consists of AI-generated paragraphs edited by MFA students, we retrained a better calibrated reward model with few fully human written high quality text (See Appendix A.11 for more details).Figure 4(a) shows writing quality scores from the WQRM model when prompts lack detailed content. Award-winning authors achieve a median score of 8.9, while LLMs score 4.8-6.6 with much higher variance. Despite WQRM being trained only on AI-generated paragraphs edited by MFA students and relatively fewer human written samples, it scored 50 author-written texts higher than all LLMs, demonstrating model generalization. GPT4.5, though considered the best writing LLM, showed no quality advantage.The significant gap between award-winning authors and LLMs shows that in the absence of original good-quality content, all LLMs are poor writers.